# FreeDrag: Feature Dragging for Reliable Point-based Image Editing

### v3

@@ 4.4 Directional Editing
$\mathcal{L}_{mask}={\left\|({F_{0}-F})\odot(1-M)\right\|_{1}},$ (10)
@@
$\mathcal{L}_{total}=\mathcal{L}_{drag}+\gamma\cdot\mathcal{L}_{drag}$ (11)
@@ 5 Experiments
Since the proposed feature dragging pipeline is constructed based on feature map, thus it can be effortlessly implemented on StyleGAN2 models [22] and latent diffusion models[40] by extracting corresponding feature maps.
@@ 5.1 Implementation Details
Parameter $r$ in Eq. 1 is set as 3, and parameter $\gamma$ in Eq. 11 is set as 10. For StyleGAN2 models, the feature map is extracted after the 6th block and the optimization for latent code is conducted in the extended $\mathcal{W}^{+}$ space[1]. We set $l=0.4$ and $d=4$ for elephant and lion models that are observed to likely perform larger movement in single optimization step, and $l=0.3$ and $d=3$ for other StyleGAN2 models. For diffusion models, following DragDiffusion[43], we fine-turn a LoRA [16] with rank of 16 on the UNet parameters for each image, which is used for both image inversion and dragging editing, and the feature map is extracted from the U-Net. We also replaces the feature map with diffusion latent in Eq. 10 to keep consistent with DragDiffusion. The parameters $l$ and $d$ are empirically set as 1 and 5 in diffusion models, respectively. To reflect the performance of different dragging pipelines themselves, FreeDrag and DragDiffusion utilize the same LoRA parameters for the the same image. To fully capture the potential of each method, the max step in set as 300 for all methods.
@@
Figure 7: Visualization of content consistency of symmetrical dragging. Reverse handle points and target points to construct a pair of symmetrical dragging instructions.
5.2 Dataset Construction
Since there is no public dataset to evaluate the drag-based editing in StyleGAN2, we propose FreeDragBench, which is the first dataset customized for GAN-based dragging editing. As presented in Table 1, FreeDragBench consists of 600 images randomly generated by five different StyleGAN2 models, equipped with 2251 dragging instructions toiled for image content (including the editing in pose, size, position, etc.), as shown in Fig. 6.
Furthermore, since the ground-truth corresponding to dragging instruction is not available, we propose a new metric to measure the accuracy of dragging editing, i.e., the Content Consistency under Symmetrical Dragging (CCSD). To be specific, as depicted Fig. 7, we reuse the reverse side of original dragging instruction to construct a symmetrical dragging instruction pair, and measure the content consistency under this symmetrical dragging instruction pair. A low CCSD value requires accurate dragging in symmetrical editing, which could be used as an effective measurement metric under the absence of ground-truth.
Figure 8: Demonstration of the edited results of FreeDrag and DragGAN in eight different StyleGAN2 models. FreeDrag advances the accuracy of point dragging and the reliablity of structure preservation.
Figure 9: Demonstration of real image editing results of FreeDrag and DragDiffusion. FreeDrag achieves more precise semantic dragging with better structure preservation.
@@ 5.2 Dataset Construction
DragGAN[33] 38.07 0.81 19.14 0.57 36.36 0.92 21.90 1.57 11.17 1.11 8.26
FreeDrag 29.50 0.32 15.67 0.21 33.50 0.49 21.18 0.83 10.86 0.76 2.74
Table 2: Quantitative evaluation on FreeDragBench. A lower FID score indicates better fidelity in single dragging editing, while lower CCSD (×$(\times 10^{2})$ scores imply higher accuracy in two symmetrical dragging editing. The time is calculated on Face category.
5.3 Qualitative Evaluation
As depicted in Fig. 8, FreeDrag successfully avoids the abnormal disappearance of handle points (e.g., the vanished eyes in human face, and the mouth of cartoon character and cat), showcasing its superiority in fine-detail editing. Meanwhile, FreeDrag achieves better stability against drastic content distortions (see the eye of horse), steadily attaining the editing intent. Moreover, FreeDrag exhibits better robustness in handling similar points, resulting in reliable and precise dragging editing, as demonstrated in the examples of third row. Additionally, FreeDrag effectively mitigates the potential misguidance during optimization steps, leading to more natural and coherent editing results, as observed in last row in Fig. 8.
For the image editing with the combination of diffusion models, FreeDrag also attains impressive performance. As shown in Fig. 9, FreeDrag outperforms DragDiffusion in both editing accuracy (see the examples from first to third columns) and structure preservation (see the examples from fourth to last columns), thus achieving superior quality of point-based dragging editing.
5.4 Quantitative Evaluation
For quantitative evaluation, we implement comparison with DragGAN and DragDiffusion in FreeDragBench and DragBench[43], respectively. Specifically, for the comparison in FreeDragBench, we use FID and the proposed CCSD to evaluate the image quality and editing accuracy, respectively. For DragBench that owns images with varying resolution, we follow the setting in DragDiffusion[43], i.e., Mean Distance (MD) for dragging accuracy measurement and LPIPS [52] for image fidelity evaluation. The mean distance is obtained by calculating the corresponding relationship of points between original image and edited image based on DIFT[45].
@@ 5.4 Quantitative Evaluation
Table 3: Quantitative evaluation on DragBench. The time consumption is computed on DragBench which only includes the dragging process because a fine-turned LoRA can be used for multiple image editing with different dragging instructions.
@@ 5.5 Ablation Study
The parameters $l$ and $d$ determines the initial feature discrepancy and maximum single movement distance, thus controls the style of total dragging editing. Specifically, a too small $l$ or $d$ implies a more conservative editing strategy, which prefers small motion and refuses large updating scale, thus failing to reach the target point in limited optimization steps, as shown in Fig. 10(b). In contrast, a too large $l$ or $d$ means a more impulsive editing strategy, which appears to accept large updating scale and larger movement distance and thus increases the risk of coarse feature updating, resulting in the damage to editing accuracy, as can be observed in Fig. 10(d).
Furthermore, we assigns $\lambda=0$ in Eq. 3 to obtain a stationary template feature to evaluate the effect of adaptive updating strategy, and adopts Eq. 7 rather than Eq. 9 to evaluate the effect of backtracking mechanism. As can be observed in Fig. 11, both of them play necessary roles for better editing quality. More examples with other variants are provided in the supplementary material.
@@ 6 Conclusion
In this work, we propose FreeDrag, a novel feature dragging framework for reliable point-based image editing. By incorporating an adaptive template feature, FreeDrag allows for flexible control in the scale of each feature updating, which contributes to stronger stability under drastic content change, resulting in a better immunity against point missing. Meanwhile, the established line search with backtracking effectively mitigates the misguidance caused by similar points and allows timely adjustment for motion plan by effectively discriminating abnormal motion, leading to reliable and continuous movements towards final target point. Extensive experiments demonstrate the reliability of FreeDrag in precise semantic dragging and stable structure preservation, indicating superior editing quality.

### v4

@@ 4.4 Directional Editing
$\mathcal{L}_{mask}={\left\|({F_{0}-F})\odot(1-M)\right\|_{1}},\vspace{-2mm}$ (10)
@@
$\mathcal{L}_{total}=\mathcal{L}_{drag}+\gamma\cdot\mathcal{L}_{mask}.\vspace{-2mm}$ (11)
@@ 5 Experiments
Since the proposed feature dragging pipeline is constructed based on the feature map, thus it can be effortlessly implemented on StyleGAN2 models [22] and latent diffusion models[40] by extracting corresponding feature maps.
@@ 5.1 Implementation Details
Parameter $r$ in Eq. 1 is set as 3, and parameter $\gamma$ in Eq. 11 is set as 10. For StyleGAN2 models, the feature map is extracted after the 6th block and the optimization for latent code is conducted in the extended $\mathcal{W}^{+}$ space[1]. We set $l=0.4$ and $d=4$ for elephant and lion models that are observed to likely perform larger movement in a single optimization step, and $l=0.3$ and $d=3$ for other StyleGAN2 models. For diffusion models, following DragDiffusion[43], we fine-tune a LoRA [16] with rank of 16 on the UNet parameters for each image, which is used for both image inversion and dragging editing, and the feature map is extracted from the U-Net. We also replace the feature map with diffusion latent in Eq. 10 to keep consistent with DragDiffusion. The parameters $l$ and $d$ are empirically set as 1 and 5 in diffusion models, respectively. To reflect the performance of different dragging pipelines themselves, FreeDrag and DragDiffusion utilize the same LoRA parameters for the same image. To fully capture the potential of each method, the max step is set as 300 for all methods.
5.2 Dataset Construction
Since there is no public dataset to evaluate the drag-based editing in StyleGAN2, we propose FreeDragBench, which is the first dataset customized for GAN-based dragging editing. As presented in Table 1, FreeDragBench consists of 600 images randomly generated by five different StyleGAN2 models, equipped with 2251 dragging instructions tailored for image content (including the editing in the pose, size, position, etc.), as shown in Fig. 6.
Furthermore, since the ground-truth corresponding to dragging instruction is not available, we propose a new metric to measure the accuracy of dragging editing, i.e., the Content Consistency under Symmetrical Dragging (CCSD). To be specific, as depicted Fig. 7, we reuse the reverse side of the original dragging instruction to construct a symmetrical dragging instruction pair and measure the content consistency under this symmetrical dragging instruction pair. To avoid penalizing stochastic elements with no effect on perception, LPIPS[52] is used for similarity measurement. A low CCSD value requires accurate dragging in symmetrical editing, which could be used as an effective measurement metric in the absence of ground-truth.
5.3 Qualitative Evaluation
As depicted in Fig. 9, FreeDrag successfully avoids the abnormal disappearance of handle points (e.g., the vanished eyes in the human face, and the mouth of cartoon character and cat), showcasing its superiority in fine-detail editing. Meanwhile, FreeDrag achieves better stability against drastic content distortions (see the eye of the horse), steadily attaining the editing intent. Moreover, FreeDrag exhibits better robustness in handling similar points, resulting in reliable and precise dragging editing, as demonstrated in the examples of the third row. Additionally, FreeDrag effectively mitigates the potential misguidance during optimization steps, leading to more natural and coherent editing results, as observed in the last row in Fig. 9.
@@ 5.3 Qualitative Evaluation
Figure 7: Visualization of the proposed CCSD metric.
Figure 8: Comparison with EditGAN[26] in editing accuracy.
Figure 9: Demonstration of the edited results of FreeDrag and DragGAN[32] in eight different StyleGAN2 models.
@@
DragGAN[33] 38.07 0.83 19.14 0.56 36.36 0.73 21.90 1.20 11.17 1.19 8.26
FreeDrag 29.50 0.35 15.67 0.23 33.50 0.37 21.18 0.68 10.86 0.82 2.74
Table 2: Quantitative evaluation on FreeDragBench. A lower FID score indicates better fidelity in single dragging editing, while lower CCSD (×10)($\times$ 10)( × 10 ) scores imply higher accuracy in two symmetrical dragging editing. The time is calculated on Face category.
@@
Table 3: Quantitative evaluation on DragBench. The time consumption is computed on DragBench which only includes the dragging process because a fine-tuned LoRA can be used for multiple image editing with different dragging instructions.
For image editing with the combination of diffusion models, FreeDrag also attains impressive performance. As shown in Fig. 10, FreeDrag outperforms DragDiffusion in both editing accuracy (see the examples from the first to third columns) and structure preservation (see the examples from the fourth to last columns), thus achieving superior quality of point-based dragging editing.
Additionally, we further conduct a comparison with EditGAN[26], which performs fine-grained editing by drawing object-level masks. As shown in Fig. 8, FreeDrag better follows editing instructions.
5.4 Quantitative Evaluation
For quantitative evaluation, we implement comparison with DragGAN and DragDiffusion in FreeDragBench and DragBench[43], respectively. Specifically, for the comparison in FreeDragBench, we use FID and the proposed CCSD to evaluate the image quality and editing accuracy, respectively. For DragBench that owns images with varying resolution, we follow the setting in DragDiffusion[43], i.e., Mean Distance (MD) for dragging accuracy measurement and LPIPS [52] for image fidelity evaluation. The mean distance is obtained by calculating the corresponding relationship of points between the original image and the edited image based on DIFT[45].
Figure 10: Demonstration of real image editing results of FreeDrag and DragDiffusion[32].
@@ 5.4 Quantitative Evaluation
Metric w/o updating w/o backtracking Ours
CCSD (×10)($\times$ 10)( × 10 ) 0.82 0.52 0.35
Table 4: Quantitative ablation on human face model.
@@ 5.5 Ablation Study
The parameters $l$ and $d$ determine the initial feature discrepancy and maximum single movement distance, thus controlling the style of total dragging editing. Specifically, a too small $l$ or $d$ implies a more conservative editing strategy, which prefers small motion and refuses large updating scale, thus failing to reach the target point in limited optimization steps, as shown in Fig. 11(b). In contrast, a too large $l$ or $d$ means a more impulsive editing strategy, which appears to accept large updating scale and larger movement distance and thus increases the risk of coarse feature updating, resulting in damage to editing accuracy, as can be observed in Fig. 11(d).
Furthermore, we assign $\lambda=0$ in Eq. 3 to obtain a stationary template feature to evaluate the effect of adaptive updating strategy and adopt Eq. 7 rather than Eq. 9 to evaluate the effect of backtracking mechanism. As can be observed in Fig. 12, both of them play necessary roles for better editing quality. The quantitative ablation in Table 4 also validates their necessity.
5.6 Limitation
Freedrag achieves image manipulation with a reliance on the 2D dragging instructions, which may be limited in certain 3D editing scenarios due to potential ambiguity. As can be observed in Fig. 13, an intention of “head up” might result in a “resized shape” instead.
Figure 13: Limitations of 2D dragging instructions. (a)-(b): Dragging instructions. (c)-(d): Edited results.
@@ 6 Conclusion
In this work, we propose FreeDrag, a novel feature dragging framework for reliable point-based image editing. By incorporating an adaptive template feature, FreeDrag allows for flexible control in the scale of each feature updating, which contributes to stronger stability under drastic content change, resulting in a better immunity against point missing. Meanwhile, the established line search with backtracking effectively mitigates the misguidance caused by similar points and allows timely adjustment for motion plan by effectively discriminating abnormal motion, leading to reliable and continuous movements towards the final target point. Extensive experiments demonstrate the reliability of FreeDrag in precise semantic dragging and stable structure preservation, indicating superior editing quality.
Acknowledgement. This work is supported in part by the Postdoctoral Fellowship Program of CPSF GZB20230713, in part by the Anhui Provincial Key Research and Development Plan 202304a05020072, in part by the Fundamental Research Funds for the Central Universities WK2090000065, and in part by the Anhui Provincial Natural Science Foundation 2308085QF226. This work is partially supported by the National Key R&D Program of China (2022ZD0160201), and Shanghai Artificial Intelligence Laboratory.